The demand and supply equations for a certain item are given by
D = –5p + 40
S = –p2 + 30p – 8
Find the equilibrium price.

Answer by stanbon(75887)

Set the demand equal to the supply and solve for "p"
D = –5p + 40
S = –p2 + 30p – 8
Find the equilibrium price.
-----------------
-5p+40 = -p^2+30p-8
p^2-35p+48 = 0
Use the quadratic formula to find "p":
p = [35+-sqrt(35^2-4*1*48]/2
p = [35 +- sqrt(1033)]/2
p = [35 +- 32.14..]/2
p = $33.57 or p = $1.43
==========
Cheers,
Stan H.




Cheers,
Stan H.